A sample of ammonia (NH3) gas is completely decomposed to nitrogen and hydrogen gases over heated iron wool. If the total pressure after the reaction is 1510mmHg, calculate the partial pressure of N2 and H2.

Grade:12

1 Answers

Anil Pannikar AskiitiansExpert-IITB
13 years ago

Dear Nitin ,

 

  2 NH3---->N2 + 3 H2

let NH3  = x mole

so, N2 = x/2 mole

H2= 3x/2 mole

mole frac. H2 = (3/2)/(1 + 3/2 + 1/2)

mole frac. N2 = (1/2)/(1 + 3/2 + 1/2)

Partial press H2 = 1510 * mole frac. H2

Partial press N2 = 1510 * mole frac. N2

 

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Anil Pannikar

IIT Bombay